OpenStudy (anonymous):

Find the general term an for the geometric sequence a1=-3 and a2=6?

OpenStudy (anonymous):

\[a_{1}=-3 and a_{2}=6\]

OpenStudy (raden):

use the formula : an = a1 r^(n-1) with a1 is given -3 and to get the ratio(r), defined r = a2/a1 = 6/-3 = -2 therefore, an = a1 * r^(n-1) an = -3 * (-2)^(n-1)